# A Generalized Difference-in-Differences Estimator for Unbiased Estimation of Desired Estimands from Staggered Adoption and Stepped-Wedge Settings

### v1

@@ Appendix A Determining the Matrix of DID Estimators
This is the matrix that, when multiplied (on the right) by the column vector of outcomes $\bm{Y}$, gives the column vector of two-by-two DID estimators $\bm{D}$. In particular, $D_{i,i^{\prime},j,j^{\prime}}$ will be found on the following row number of the vector $\bm{D}$:
@@
Each row $k$ can then be associated with the corresponding $D_{i,i^{\prime},j,j^{\prime}}$. From the values of $T_{i}$ and $T_{i^{\prime}}$, along with $j$ and $j^{\prime}$, the type (from 1–6) of this two-by-two DID estimator can then be determined. And its expectation under the desired assumption setting can be determined as well. Thus, the $\binom{J}{2}\binom{N}{2}\times 1$ column vector $E[\bm{D}]$ and associated matrix $\bm{F}$ can be constructed in terms of the estimands possible for the assumption setting.
@@ B.1 Lemmata for Proof of Theorem 3
Lemma 4 Let $\bm{F}$ and $\bm{A}$ be as defined previously. Then $ker(\bm{A}^{T})\subset ker(\bm{F}^{T})$.
@@
Lemma 5 Let $\bm{A}$ be as defined previously for a setting with $N\geq 2$ units and $J\geq 2$ periods. Then $rank(A)=rank(A^{T})=(N-1)(J-1)$.
@@ B.2 Proof of Theorem 3
Theorem 3 Let $\bm{F}$, $\bm{w}$, $\bm{v}$, and $\bm{d}$ be as defined previously. Then the following are true about the set of estimators of the form $\hat{\theta}=\bm{w}^{T}\bm{d}$ that are unbiased for $\theta_{e}$ under the assumption setting:
@@
Let $\bm{w}_{1}\in W$ be a solution to $\bm{F}^{T}\bm{w}=\bm{v}$. Any other solution to $\bm{F}^{T}\bm{w}=\bm{v}$ can be expressed as $\bm{w}_{1}+\bm{x}$, where $\bm{x}\in ker(\bm{F}^{T})$. The vector space of $ker(\bm{F}^{T})$ has dimension $nullity(\bm{F}^{T})=\binom{N}{2}\binom{J}{2}-rank(\bm{F}^{T})$, corresponding to the dimension of the affine space $W$ as found above. By Lemma 4, $ker(\bm{A}^{T})\subset ker(\bm{A}^{T})$, and so $ker(\bm{F}^{T})$ can be written as the direct sum of $ker(\bm{A}^{T})+K$, where $K$ is defined as the orthogonal complement of $ker(\bm{A}^{T})$ within $ker(\bm{F}^{T})$. So we can further express $W=\{\bm{w}_{1}+\bm{x}_{1}+\bm{x}_{2}:~{}\bm{x}_{1}\in ker(\bm{A}^{T}),~{}\bm{x}_{2}\in K\}$. However, $\bm{A}^{T}\bm{x}_{1}=\bm{0}$ if $\bm{x}_{1}\in ker(\bm{A}^{T})$. Hence, for any $\bm{w}\in W$ and any $\bm{x}_{1}\in ker(\bm{A}^{T})$, $\bm{w}^{T}\bm{A}=(\bm{w}+\bm{x}_{1})^{T}\bm{A}$, and so the estimators $\hat{\theta}_{1}$ and $\hat{\theta}_{2}$ defined by the weight vectors $\bm{w}$ and $\bm{w}+\bm{x}_{1}$ are equal for all $\bm{y}$. Thus, the only unique estimators (in terms of the underlying observations) are given by the subspace $W_{A}=\{\bm{w}_{1}+\bm{x}_{2}:~{}\bm{x}_{w}\in K\}$.
@@
This can be further simplified using the rank-nullity theorem (see George & Ajayakumar (2024), p. 89) and Lemma 5:
@@
Appendix C Supplementary Results
@@
S5 $\hat{\theta}^{(5)}$ 0.771 0.014 0.771 0.009 0.772 0.012
S4 $\sum_{j=2}^{7}\hat{\theta}^{(4)}_{j}$ 0.784 0.024 0.784 0.031 0.785 0.029
S3 $\sum_{a=1}^{7}\hat{\theta}^{(3)}_{a}$ 0.828 0.391 0.828 0.397 0.827 0.364
S2 $\sum_{j=2}^{7}\sum_{a=1}^{j-1}\hat{\theta}^{(2)}_{j,a}$ 0.801 0.168 0.801 0.160 0.802 0.160
@@
S5 $\hat{\theta}^{(5)}$7.35% 0.020 -7.35% 0.015 -7.33% 0.024
S4 $\sum_{j=2}^{7}\hat{\theta}^{(4)}_{j}$6.81% 0.039 -6.81% 0.052 -6.78% 0.045
S3 $\sum_{a=1}^{7}\hat{\theta}^{(3)}_{a}$5.57% 0.430 -5.57% 0.423 -5.58% 0.411
S2 $\sum_{j=2}^{7}\sum_{a=1}^{j-1}\hat{\theta}^{(2)}_{j,a}$6.11% 0.240 -6.11% 0.243 -6.09% 0.230

### v3

@@ Appendix A Determining the Matrix of DID Estimators
This is the matrix that, when multiplied (on the right) by the column vector of outcomes $\bm{Y}$, gives the column vector of two-by-two DID estimators $\bm{d}$, defined in Section 2.4. In particular, $D_{i,i^{\prime},j,j^{\prime}}$ will be found on the following row number of the vector $\bm{d}$:
@@
Each row $k$ can then be associated with the corresponding $D_{i,i^{\prime},j,j^{\prime}}$. From the values of $T_{i}$ and $T_{i^{\prime}}$, along with $j$ and $j^{\prime}$, the type (from 1–6) of this two-by-two DID estimator can then be determined using Table 1. And its expectation under the specified assumption setting can be determined as well using Table 3. Thus, the $\binom{J}{2}\binom{N}{2}\times 1$ column vector $E[\bm{d}]$ and associated matrix $\bm{F}$ can be constructed in terms of the estimands possible for the assumption setting.
@@ B.1 Lemmata for Proof of Theorem 3
Lemma 3.1 Let $\bm{F}$ and $\bm{A}$ be as defined in Section 2. Then $ker(\bm{A}^{T})\subset ker(\bm{F}^{T})$.
@@
Lemma 3.2 Let $\bm{A}$ be as defined in Section 2 for a setting with $N\geq 2$ units and $J\geq 2$ periods. Then $rank(\bm{A})=rank(\bm{A}^{T})=(N-1)(J-1)$.
@@ B.2 Proof of Theorem 3
Theorem 3 Let $\bm{F}$, $\bm{w}$, $\bm{v}$, and $\bm{d}$ be as defined in Section 2. Then the following are true about the set of estimators of the form $\hat{\theta}=\bm{w}^{T}\bm{d}$ that are unbiased for $\theta_{e}$ under the assumption setting:
@@
Let $\bm{w}_{1}\in W$ be a solution to $\bm{F}^{T}\bm{w}=\bm{v}$. Any other solution to $\bm{F}^{T}\bm{w}=\bm{v}$ can be expressed as $\bm{w}_{1}+\bm{x}$, where $\bm{x}\in ker(\bm{F}^{T})$. The vector space of $ker(\bm{F}^{T})$ has dimension $nullity(\bm{F}^{T})=\binom{N}{2}\binom{J}{2}-rank(\bm{F}^{T})$, corresponding to the dimension of the affine space $W$ as found above. By Lemma 3.1, $ker(\bm{A}^{T})\subset ker(\bm{A}^{T})$, and so $ker(\bm{F}^{T})$ can be written as the direct sum of $ker(\bm{A}^{T})+K$, where $K$ is defined as the orthogonal complement of $ker(\bm{A}^{T})$ within $ker(\bm{F}^{T})$. So we can further express $W=\{\bm{w}_{1}+\bm{x}_{1}+\bm{x}_{2}:~{}\bm{x}_{1}\in ker(\bm{A}^{T}),~{}\bm{x}_{2}\in K\}$. However, $\bm{A}^{T}\bm{x}_{1}=\bm{0}$ if $\bm{x}_{1}\in ker(\bm{A}^{T})$. Hence, for any $\bm{w}\in W$ and any $\bm{x}_{1}\in ker(\bm{A}^{T})$, $\bm{w}^{T}\bm{A}=(\bm{w}+\bm{x}_{1})^{T}\bm{A}$, and so the estimators $\hat{\theta}_{1}$ and $\hat{\theta}_{2}$ defined by the weight vectors $\bm{w}$ and $\bm{w}+\bm{x}_{1}$ are equal for all $\bm{y}$. Thus, the only unique estimators (in terms of the underlying observations) are given by the subspace $W_{A}=\{\bm{w}_{1}+\bm{x}_{2}:~{}\bm{x}_{w}\in K\}$.
@@
This can be further simplified using the rank-nullity theorem (see George & Ajayakumar (2024), p. 89) and Lemma 3.2:
@@
Appendix C Details of Simulation Estimators
The overall effect estimators used in the simulations are detailed in Table 9.
Table 9: Estimators targeting overall effects used for comparisons in simulations. Number corresponds to the x-axis in Figs. 3 and 4. Type corresponds to the labels in those figures, with GD representing a generalized DID estimator, SA a staggered adoption estimator (see Section 4), and ME a mixed effects model estimator. Assumption corresponds to the effect heterogeneity assumption setting outlined in Table 2; for non-GD estimators, the assumption setting is inferred from the model and estimator. Unless otherwise noted: $a_{i}$ and $b_{j}$ refer to fixed effects for cluster $i$ and period $j$, respectively; $\alpha_{i}$, $\nu_{ij}$, and $\epsilon_{ijk}$ refer to random effects for cluster, cluster-period, and individual respectively and follow independent normal distributions with mean 0 and variances $\sigma^{2}_{\alpha}$, $\sigma^{2}_{\nu}$, and $\sigma^{2}_{e}$, respectively; and $X_{ij}$ is the binary treatment indicator for cluster $i$ in period $j$ with $T_{i}$ equal to the first period for cluster $i$ where $X_{ij}=1$. Estimator for GD estimators uses the notation from this paper; for ME estimators, it uses the defined model; and for other estimators, it follows the cited source as closely as possible.
No. Type Assumption Model Estimator
Overall Estimand
1 GD S5 Independent Covariance $\hat{\theta}^{(5)}$
2 SA S5 TW: two-way fixed effects
$E[Y_{ijk}]=a_{i}+b_{j}+\theta X_{ij}$
$\hat{\theta}$
3 ME S5 $Y_{ijk}=\mu+\alpha_{i}+b_{j}+\nu_{ij}+\theta X_{ij}+\epsilon_{ijk}$ $\hat{\theta}$
4 SA S5 CO-3: equal weights
(Kennedy‐Shaffer et al. (2020))
$\hat{\theta}^{CO-3}$
Calendar-Time Averaged Estimand
5 GD S4 Independent Covariance $\frac{1}{6}\sum_{j=2}^{7}\hat{\theta}^{(4)}_{j}$
6 ME S4 $Y_{ijk}=\mu+\alpha_{i}+b_{j}+\nu_{ij}+\theta_{j}X_{ij}+\epsilon_{ijk}$ $\frac{1}{6}\sum_{j=2}^{7}\hat{\theta}_{j}$
7 SA S4 CS: calendar ATT
(Callaway & Sant’Anna (2021))
$\frac{1}{6}\sum_{t=2}^{7}\hat{\theta}_{c}(t)$
8 GD S2 Independent Covariance $\frac{1}{6}\sum_{j=2}^{7}\left[\frac{1}{j-1}\sum_{a=1}^{j-1}\hat{\theta}^{(2)}_{j,a}\right]$
9 ME S2 $Y_{ijk}=\mu+\alpha_{i}+b_{j}+\nu_{ij}+\theta_{j,a}X_{ij}I(T_{i}-j+1=a)+\epsilon_{ijk}$ $\frac{1}{6}\sum_{j=2}^{7}\left[\frac{1}{j-1}\sum_{a=1}^{j-1}\hat{\theta}_{j,a}\right]$
No. Type Assumption Model Estimator
Exposure-Time Averaged Estimand
10 GD S3 Independent Covariance $\frac{1}{7}\sum_{a=1}^{7}\hat{\theta}^{(3)}_{a}$
11 ME S3 $Y_{ijk}=\mu+\alpha_{i}+b_{j}+\nu_{ij}+\theta_{a}X_{ij}I(T_{i}-j+1=a)+\epsilon_{ijk}$ $\frac{1}{7}\sum_{a=1}^{7}\hat{\theta}_{a}$
12 GD S3 Independent Covariance $\frac{1}{6}\sum_{a=1}^{6}\hat{\theta}^{(3)}_{a}$
13 SA S3 CS: dynamic ATT
(Callaway & Sant’Anna (2021))
$\frac{1}{6}\sum_{e=1}^{6}\hat{\theta}_{es}(e)$
14 GD S2 Independent Covariance $\frac{1}{6}\sum_{a=1}^{6}\left[\frac{1}{6-a}\sum_{j=a+1}^{7}\hat{\theta}^{(2)}_{j,a}\right]$
15 ME S2 $Y_{ijk}=\mu+\alpha_{i}+b_{j}+\nu_{ij}+\theta_{j,a}X_{ij}I(T_{i}-j+1=a)+\epsilon_{ijk}$ $\frac{1}{6}\sum_{a=1}^{6}\left[\frac{1}{6-a}\sum_{j=a+1}^{7}\hat{\theta}_{j,a}\right]$
Group Averaged Estimand
16 GD S2 Independent Covariance $\frac{1}{6}\sum_{a=1}^{6}\sum_{j=a+1}^{7}\frac{1}{7-j+a}\hat{\theta}_{j,a}^{(2)}$
17 SA S2 CS: group ATT
(Callaway & Sant’Anna (2021))
$\hat{\theta}^{O}_{sel}=\frac{1}{6}\sum_{\tilde{g}=2}^{7}\hat{\theta}_{sel}(\tilde{g})$
Average Treatment Effect on the Treated (Cluster-Period Averaged) Estimand
18 GD S2 Independent Covariance $\frac{1}{21}\sum_{j=2}^{7}\sum_{a=1}^{j-1}\hat{\theta}_{j,a}^{(2)}$
19 SA S2 CS: simple ATT
(Callaway & Sant’Anna (2021))
$\hat{\theta}^{O}_{W}$
20 SA S2 SA: Interaction-Weighted ATT
(Sun & Abraham (2021))
$\frac{1}{6}\sum_{g=2}^{7}\hat{v}_{g}$
21 ME S2 $Y_{ijk}=\mu+\alpha_{i}+b_{j}+\nu_{ij}+\theta_{j,a}X_{ij}I(T_{i}-j+1=a)+\epsilon_{ijk}$ $\frac{1}{21}\sum_{a=1}^{6}\sum_{j=a+1}^{7}\hat{\theta}_{j,a}$
The period-specific effect estimators used in the simulations are detailed in Table 10.
Table 10: Estimators targeting period-specific effects used for comparisons in simulations. Number corresponds to the x-axis in Figs. 7 and 8. Type corresponds to the labels in those figures, with GD representing a generalized DID estimator, SA a staggered adoption estimator (see Section 4), and ME a mixed effects model estimator. Assumption corresponds to the effect heterogeneity assumption setting outlined in Table 2; for non-GD estimators, the assumption setting is inferred from the model and estimator. Unless otherwise noted: $a_{i}$ and $b_{j}$ refer to fixed effects for cluster $i$ and period $j$, respectively; $\alpha_{i}$, $\nu_{ij}$, and $\epsilon_{ijk}$ refer to random effects for cluster, cluster-period, and individual respectively and follow independent normal distributions with mean 0 and variances $\sigma^{2}_{\alpha}$, $\sigma^{2}_{\nu}$, and $\sigma^{2}_{e}$, respectively; and $X_{ij}$ is the binary treatment indicator for cluster $i$ in period $j$ with $T_{i}$ equal to the first period for cluster $i$ where $X_{ij}=1$. Estimator for GD estimators uses the notation from this paper; for ME estimators, it uses the defined model; and for other estimators, it follows the cited source as closely as possible.
No. Type Assumption Model Estimator
Calendar Period 3 Estimand
1 GD S4 Independent Covariance $\hat{\theta}^{(4)}_{3}$
2 ME S4 $Y_{ijk}=\mu+\alpha_{i}+b_{j}+\nu_{ij}+\theta_{j}X_{ij}+\epsilon_{ijk}$ $\hat{\theta}_{3}$
3 GD S2 Independent Covariance $\frac{1}{2}\sum_{a=1}^{2}\hat{\theta}^{(2)}_{3,a}$
4 ME S2 $Y_{ijk}=\mu+\alpha_{i}+b_{j}+\nu_{ij}+\theta_{j,a}X_{ij}I(T_{i}-j+1=a)+\epsilon_{ijk}$ $\frac{1}{2}\sum_{a=1}^{2}\hat{\theta}_{3,a}$
Exposure Period 2 Estimand
5 GD S3 Independent Covariance $\hat{\theta}^{(3)}_{2}$
6 ME S3 $Y_{ijk}=\mu+\alpha_{i}+b_{j}+\nu_{ij}+\theta_{a}X_{ij}I(T_{i}-j+1=a)+\epsilon_{ijk}$ $\hat{\theta}_{2}$
7 GD S2 Independent Covariance $\frac{1}{5}\sum_{j=3}^{7}\hat{\theta}^{(2)}_{j,2}$
8 ME S2 $Y_{ijk}=\mu+\alpha_{i}+b_{j}+\nu_{ij}+\theta_{j,a}X_{ij}I(T_{i}-j+1=a)+\epsilon_{ijk}$ $\frac{1}{6}\sum_{j=3}^{8}\hat{\theta}_{j,2}$
No. Type Assumption Model Estimator
Exposure Period 1 Estimand
9 GD S3 Independent Covariance $\hat{\theta}^{(3)}_{1}$
10 ME S3 $Y_{ijk}=\mu+\alpha_{i}+b_{j}+\nu_{ij}+\theta_{a}X_{ij}I(T_{i}-j+1=a)+\epsilon_{ijk}$ $\hat{\theta}_{1}$
11 SA S2 CO-2: proportional weights
(Kennedy‐Shaffer et al. (2020))
$\hat{\theta}^{CO-2}$
12 GD S2 Independent Covariance $\frac{1}{6}\sum_{j=2}^{7}\hat{\theta}^{(2)}_{j,1}$
13 ME S2 $Y_{ijk}=\mu+\alpha_{i}+b_{j}+\nu_{ij}+\theta_{j,a}X_{ij}I(T_{i}-j+1=a)+\epsilon_{ijk}$ $\frac{1}{6}\sum_{j=2}^{8}\hat{\theta}_{j,1}$
14 SA S2 CH: first-difference estimator
(de Chaisemartin & D’Haultfœuille (2020))
$\hat{DID}_{M}$
15 SA S2 CO-1: equal weights
(Kennedy‐Shaffer et al. (2020))
$\hat{\theta}^{CO-1}$
Appendix D Supplementary Results
@@
S5 $\hat{\theta}^{(5)}$ 0.771 0.008 0.771 0.013 0.772 0.007
S4 $\frac{1}{6}\sum_{j=2}^{7}\hat{\theta}^{(4)}_{j}$ 0.784 0.026 0.784 0.021 0.785 0.029
S3 $\frac{1}{7}\sum_{a=1}^{7}\hat{\theta}^{(3)}_{a}$ 0.828 0.371 0.828 0.413 0.827 0.386
S2 $\frac{1}{21}\sum_{j=2}^{7}\sum_{a=1}^{j-1}\hat{\theta}^{(2)}_{j,a}$ 0.801 0.155 0.801 0.149 0.802 0.162
@@
S5 $\hat{\theta}^{(5)}$7.35% 0.013 -7.35% 0.020 -7.33% 0.0013
S4 $\frac{1}{6}\sum_{j=2}^{7}\hat{\theta}^{(4)}_{j}$6.81% 0.054 -6.81% 0.054 -6.78% 0.048
S3 $\frac{1}{7}\sum_{a=1}^{7}\hat{\theta}^{(3)}_{a}$5.57% 0.409 -5.57% 0.456 -5.58% 0.430
S2 $\frac{1}{21}\sum_{j=2}^{7}\sum_{a=1}^{j-1}\hat{\theta}^{(2)}_{j,a}$6.11% 0.220 -6.11% 0.229 -6.09% 0.232
Figure 5: Weights on cluster-period observations for the estimator $\frac{1}{6}\sum_{j=2}^{7}\hat{\theta}^{(4)}_{j}$ fitted under an exchangeable variance structure with $\rho=0.003$.
Table 13: Cumulative percentage difference estimates and permutation test p-values for generalized DID estimators of summary effects of vaccine lotteries on COVID-19 vaccine uptake, United States Midwest region, May–July 2021. Estimation occurred using an independent working correlation structure with parameter $\rho=0.95$.
Description Estimator Estimate P-Value
Overall ATT $\frac{1}{26}\sum_{\{j,a\}}\hat{\theta}_{j,a}^{(2)}$ 1.318 0.265
First-Period Effect $\frac{1}{4}\sum_{j}\hat{\theta}_{j,1}^{(2)}$ 1.311 0.044
Second-Period Effect $\frac{1}{4}\sum_{j}\hat{\theta}_{j,2}^{(2)}$ 1.570 0.027
Average Four-Week Effect $\frac{1}{12}\sum_{\{j,a:~{}j-a\leq 26,1\leq a\leq 4\}}\hat{\theta}_{j,a}^{(2)}$ 1.424 0.136
Average 2–4-Week Effect $\frac{1}{9}\sum_{\{j,a:~{}j-a\leq 26,2\leq a\leq 4\}}\hat{\theta}_{j,a}^{(2)}$ 1.477 0.143
State-Averaged Effect $\frac{1}{4}\sum_{t=15}^{30}\frac{1}{31-t}\sum_{a=1}^{31-t}\hat{\theta}_{a+t,a}^{(2)}$ 1.593 0.073
Ohio Effect $\frac{1}{12}\sum_{a=1}^{12}\hat{\theta}_{a+18,a}^{(2)}$0.016 0.987
Illinois Effect $\frac{1}{7}\sum_{a=1}^{7}\hat{\theta}_{a+23,a}^{(2)}$ 4.010 0.027
Figure 6: Weights on cluster-period observations for the overall ATT estimator (top) and average 2–4-week effect estimator (bottom) under assumption S2 fitted under an independent working correlation structure.
Figure 7: Simulation results (mean estimate plus or minus one standard deviation of estimates across simulations) for 1,000 simulations each, by simulation scenario and estimator for period-specific estimators. Dotted horizontal lines represent the true value of the target estimand.
Figure 8: Simulation results (empirical power at $\alpha=0.05$ level) for 1,000 simulations each, by simulation scenario and estimator for period-specific estimators. Note that biased estimators or those targeting substantially different estimands are excluded.